OpenStudy (missmeow):

A standard showerhead in Reyna's house dispenses 9 gallons of water per minute. Reyna changed this showerhead to an energy-saving one. The equation shows the amount of water dispensed, y, as a function of the number of minutes, x, for the new showerhead. y = 3x How much water does Reyna save each day with the change in showerheads if she uses the shower for 8 minutes each day? A) 14 gallons B) 48 gallons C) 69 gallons D) 96 gallons (Could you please also explain to me how you got to that conclusion?)

OpenStudy (donnie1999):

c

OpenStudy (xmissalycatx):

@donnie1999 Don't give direct answers please!

OpenStudy (missmeow):

how did you get there though?

OpenStudy (missmeow):

oops I meant how did you come up with your conclusion?

OpenStudy (mathstudent55):

Here is the explanation. Her old showerhead used 9 gallons per minute. She takes an 8-minute shower. How much water did she use wih the old showerhead?

OpenStudy (missmeow):

72

OpenStudy (mathstudent55):

Correct.

OpenStudy (xmissalycatx):

Okay so what you will want to do is multiply your initial amount of gallons that she used by 8, and then multiply her new amount of gallons used by 8. You would need to subtract that from the initial.

OpenStudy (mathstudent55):

Now we need to see how much water the new showerhead uses.

OpenStudy (mathstudent55):

The new showerhead uses y = 3x gallons of water. y is the number of gallons used, and x is the number of minutes the shower has been running. If she runs the shower for 1 minute, she uses y = 3 * 1 = 3 3 gallons of water. If she runs the shower 5 minutes, she uses y = 3 * 5 = 15 gallons of water.

OpenStudy (mathstudent55):

We are told she takes an 8-minute shower. How much water does the new showerhead use in 8 minutes? y = 3x = 3 * 8 = ?

OpenStudy (missmeow):

24

OpenStudy (missmeow):

?

OpenStudy (mathstudent55):

Correct. She used to use 72 gallons of water per shower. Now she uses 24 gallons of water per shower. How much water does she save with the new showerhead?

OpenStudy (missmeow):

48 gallons of water. so the answers B

OpenStudy (xmissalycatx):

Yes good job! :)

OpenStudy (missmeow):

Yay! Thanks!

OpenStudy (xmissalycatx):

Good job @mathstudent55 :3

OpenStudy (mathstudent55):

@MissMeow Correct. The answer is 72 gallons - 24 gallons = 48 gallons. As you can see the answer is not C. You did well in solving the problem, and you were smart to ask the person who answered C to explain.

OpenStudy (mathstudent55):

Thanks.